simplify the following

  1. 32343-5
  2. (70/72)-1
  3. (2x3y4/xy7)2
  4. (-2x2y3)3
  5. (81/100)-3/2
  6. (-8)2/3
  7. (x-1-y-1)/(x-y)-1
1
Expert's answer
2021-03-16T07:10:28-0400

1.

"3^23^43^{-5}=3^{2+4-5}=3"

2.

"(7^0\/7^2)^{-1}=7^{-2(-1)}=49"

3.

"(2x^3y^4/xy^7)^2=(\dfrac{2x^2}{y^3})^2=\dfrac{4x^4}{y^6}"

4.

"(-2x^2y^3)^3=-8x^6y^9"

5.

"(81/100)^{-3/2}=\dfrac{(100^{1/2})^3}{(81^{1/2})^3}=\dfrac{1000}{729}"

6.

"(-8)^{2\/3}=(64)^{1\/3}=4"

7.

"(x^{-1}-y^{-1})/(x-y)^{-1}=(\dfrac{1}{x}-\dfrac{1}{y})(x-y)"

"=-\dfrac{(x-y)^2}{xy}"